OpenStudy (anonymous):

Find the number of sides of a regular polygon if one interior angle is 120 degrees. 3 4 5 6

OpenStudy (jdoe0001):

\(\bf \textit{sum of interior angles of a regular polygon}=n\theta=180(n-2) \\ \quad \\ n=\textit{number of sides} \\ \quad \\ \theta=120^o\qquad n(120)=180(n-2)\qquad \textit{solve for "n"}\)